OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. E(-2, 2), F(5, 1)

OpenStudy (anonymous):

slope m=(2-1)/(-2-5)=1/-7=-1/7 eq of the line y-y1=m(x-x1) lets use (-2,2) y-2=(-1/7)(x-(-2))=(-1(x+2)/7) y=((-x-2)/7) +2 y=(-x/7) -(2/7)+2 y= (-x/7) +(12/7) ans